A manufacturer produces televisions. She estimates that by selling them for x dollars each, she will be able to sell 105−2x each week.

Step 2 of 2 :  

Determine what the maximum revenue will be. (Round your answer to two decimal places.)
Answer #1

& Solution :- Selling price of Each televisions = $x. No. of televisions sold = 105-22 Revenue function : R = x (205-22) R =